A spherical balloon with radius "r" has a volume as shown. Find a function that represents the amount of air required to inflate the balloon from a radius of "r" inches to a radius of (r+1) inches.

V(r)=4/3pi(r3)


1
Expert's answer
2021-10-27T12:57:44-0400

For "r" in inches and "V(r)" in inches3

"V(r) =\dfrac{4}{3}\pi r^3\\\ \\V(r+1) =\dfrac{4}{3}\pi (r+1)^3"


The amount of air to add is the difference between V(r) and V(r+1). Subtract V(r):

"V(r+1)-V(r)=\dfrac{4}{3}\pi (r+1)^3-\dfrac{4}{3}\pi (r)^3"


"=\dfrac{4}{3}\pi[(r+1)^3-(r)^3]\\\ \\= \dfrac{4}{3}\pi[r^3+3r^2+3r+1-r^3]\\\ \\=\dfrac{4}{3}\pi [3r^2+3r+1]""\ \ inches^3"



This is the final function.